JP launches a rocket straight upwards from a height of 20 feet above the ground with an initial upward velocity of 85 feet per second. The equation h(t) = -16t² + 85t + 20 models the height of the rocket off of the ground, t seconds after the rocket is launched.

a) Find the height of the rocket 3 seconds after it is launched.
b) At what time will the rocket reach the maximum height? What is the maximum height of the rocket?
c) When will the rocket reach a height of 100 feet?
d) At what time will the rocket hit the ground?

Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52788)

.
See the lessons
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower


Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Answer by addingup(3677)

h(t) = -16t² + 85t + 20
when t= 3:
h(3) = -16(3)² + 85(3) + 20 = -16(9)+ 255 + 20 = 131 feet is the height of the rocket after 3 seconds.